# Learning Getting-Up Policies for Real-World Humanoid Robots

### v1

@@ Abstract
Automatic fall recovery is a crucial prerequisite before humanoid robots can be reliably deployed. Hand-designing controllers for getting up is difficult because of the varied configurations a humanoid can end up in after a fall and the challenging terrains humanoid robots are expected to operate on. This paper develops a learning framework to produce controllers that enable humanoid robots to get up from varying configurations on varying terrains. Unlike previous successful applications of humanoid locomotion learning, the getting-up task involves complex contact patterns, which necessitates accurately modeling the collision geometry and sparser rewards. We address these challenges through a two-phase approach that follows a curriculum. The first stage focuses on discovering a good getting-up trajectory under minimal constraints on smoothness or speed / torque limits. The second stage then refines the discovered motions into deployable (i.e. smooth and slow) motions that are robust to variations in initial configuration and terrains. We find these innovations enable a real-world G1 humanoid robot to get up from two main situations that we considered: a) lying face up and b) lying face down, both tested on flat, deformable, slippery surfaces and slopes (e.g., sloppy grass and snowfield). To the best of our knowledge, this is the first successful demonstration of learned getting-up policies for human-sized humanoid robots in the real world. Project page: https://humanoid-getup.github.io/
Figure 1: HumanUP provides a simple and general two-stage training method for humanoid getting-up tasks, which can be directly deployed on Unitree G1 humanoid robots [70]. Our policies showcase robust and smooth behavior that can get up from diverse lying postures (both supine and prone) on varied terrains such as grass slopes and stone tile.
@@ I Introduction
This paper develops learned controllers that enable a humanoid robot to get up from varied fall configurations on varied terrains. Humanoid robots are susceptible to falls, and their reliance on humans for fall recovery hinders their deployment. Furthermore, as humanoid robots are expected to work in environments involving complex terrains and tight workspaces (i.e. challenging scenarios that are too difficult for wheeled robots), a humanoid robot may end up in an unpredictable configuration upon a fall, or may be on an unknown terrain. 26 of the 46 trials at the DARPA Robotics Challenge (DRC) had a fall and 25 of these falls required human intervention for recovery [40]. The DRC identified fall prevention and recovery as a major topic needing more research. This paper pursues this research and proposes a learning-based framework for learning fall recovery policies for humanoid robots under varying conditions.
@@
We present HumanUP, which circumvents these problems via a two-stage reinforcement learning (RL) based training. Stage I targets task solving in easier settings (sparse task rewards with weak regularization), while Stage II solves the task of making the learned motion deployable (i.e., control should be smooth; velocities and executed torques should be small; etc). To be specific, Stage I is designed for motion discovery without being limited by smoothness in motion or speed / torque limits, which tackles a hard task under sparse and under-specified task rewards while being less constrained for ensuring task solving. Stage II is optimized to track the state trajectory discovered in the first stage to tackle easier motion tracking with dense tracking rewards, which is under strict Sim2Real control regularization for ensuring Sim2Real transfer. From Stage I to Stage II, we employ a Sim2Real learning curriculum that progresses from simplified $\rightarrow$ full collision mesh, canonical $\rightarrow$ random initial lying posture, and weak to strong control regularization and domain randomization. This two-stage approach integrates a hard-to-easy task-solving curriculum with an easy-to-hard Sim2Real curriculum, both of which are crucial for successful learning, as demonstrated in our experiments.
We conduct experiments in simulation and the real world with the G1 platform from Unitree. In the real world, we find our framework enables the G1 robot to get up from three situations that we considered: a) supine (lying face up) poses and b) prone (lying face down) poses. This expands the capability that the G1 comes with: the default hand-crafted getting-up controller only successfully gets up from supine poses on a flat surface without bumps. In simulated experiments, we find that our framework can successfully learn getting-up policies that work on varied terrains, varied starting poses.
@@ II-A Humanoid Control
Controlling a high degree of freedom humanoid robots have fascinated researchers for the last several decades. Model-based techniques, such as those based on the Zero Moment Point (ZMP) principle [71, 31, 61, 69], optimization [41, 4, 14], and Model Predictive Control (MPC) [12, 21, 15, 74], have demonstrated remarkable success in fundamental locomotion tasks like walking, running and jumping. However, these approaches often struggle to generalize or adapt to novel environments. In contrast, learning-based approaches have recently made significant strides, continuously expanding the generalization capabilities of humanoid locomotion controllers.
Figure 2: HumanUP system overview. Our getting-up policy (Section III-A) is trained in simulation using two-stage RL training, after which it is directly deployed in the real world. (a) Stage I (Section III-B1) learns a discovery policy $f$ that figures out a getting-up trajectory with minimal deployment constraints. (b) Stage II (Section III-B2) converts the trajectory discovered by Stage I into a policy $\pi$ that is deployable, robust, and generalizable. This policy $\pi$ is trained by learning to track a slowed down version of the discovered trajectory under strong control regularization on varied terrains and from varied initial poses. (c) The two-stage training induces a curriculum (Section III-C). Stage I targets motion discovery in easier settings (simpler collision geometry, same starting poses, weak regularization, no variations in terrain), while Stage II solves the task of making the learned motion deployable and generalizable.
@@ II-B1 Getting up via motion planning
Early work from Morimoto and Doya [54] solved the getting-up problem for a two-joint, three-link walking robot in 2D, and several discrete states are used as subgoals to transit via hierarchical RL. This line of work can be viewed as an application of motion planning by configuration graph transition learning [39], where stored robot states between lying and standing are used as graph nodes to transit [20, 36, 37, 64]. More recently, some progress has been made to enable toy-sized humanoid robots to get up. For example, González-Fierro et al. [23] explores getting up from a canonical sitting posture with motion planning by imitating human demonstration with ZMP ceriterion. To address the high-dimensionality of humanoid configurations, Jeong and Lee [33] leverage bilateral symmetry to reduce the control DoFs by half and a clustering technique is used for further reducing the complexity of configuration space, thereby improving getting-up learning efficiency. However, such state machine learning using predefined configuration graphs may not be sufficient for generalizing to unpredictable initial and intermediate states, which happens when the robot operates on challenging terrains.
@@ II-B2 Hand-designed getting-up trajectories
Another solution, often adopted by commercial products, is to replay a manually designed motion trajectory. For example, Unitree [70] has a getting-up controller built into G1’s default controllers. Booster Robotics [1] designed a specific recovery controller for their robots that can help the robot recover from fallen states. The main drawback of such pre-defined trajectory getting-up controllers is that they can only handle a limited number of fallen states and lack generalization, as our experimental comparisons will show.
@@ II-B3 Learned getting-up policies for real robots
RL followed by sim-to-real has also been successfully applied for quadruped [43, 35, 72, 51] fall recovery. For example, Lee et al. [43] explore sim2real RL to achieve real-world quadruped fall recovery from complex configurations. Ji et al. [35] train a recovery policy that enables the quadruped to dribble in snowy and rough terrains continuously. Wang et al. [72] develop a quadruped recovery policy in highly dynamic scenarios..
@@ II-B4 Learned getting-up policies for character animation
A parallel research effort in character animation, also explores the design of RL-based motion imitation algorithms: DeepMimic [55], AMP [56], PHC [49], among others [11, 50, 67, 5, 22, 73]. These have also demonstrated successful getting-up controllers in simulation. By tracking user-specified getting-up curves, Frezzato et al. [17] enable humanoid characters to get up by synthesizing physically plausible motion. Without recourse to mocap data, such naturalistic getting-up controllers for simulated humanoid characters can also be developed with careful curriculum designs [65]. Some works explore sampling-based methods for addressing contact-rich character locomotion including getting up [45, 28, 57], while some works have demonstrated success in humanoid getting up with online model-predictive control [66]. It is worth noticing, however, that these works use humanoid characters with larger DoFs compared to humanoid robots (e.g., 69 DoFs in SMPL [47]) and use simplified dynamics. As a result, learned policies operate body parts at high velocities and in infeasible ways, leading to behavior that cannot be transferred into the real world directly. Hence, developing generalizable recovery controllers for humanoid robots remains an open problem.
@@ III HumanUP: Sim-to-Real Humanoid Getting Up
Our goal is to learn a getting-up policy $\pi$ that enables a humanoid to get up from arbitrary initial postures. We consider getting up from two families of lying postures: a) supine poses (i.e. lying facing up) and b) prone poses (i.e. lying face down). Getting up from these two groups of postures may require different behaviors, which makes it challenging to learn a single policy that handles both scenarios. To tackle this issue, we decompose the getting-up task from a prone pose to first rolling over and then standing up from the resulting supine posture. Therefore, our objective is to learn policies for rolling over from a prone pose and getting up from a supine pose separately.
To solve these two tasks, we propose HumanUP, a general learning framework for training getting-up and rolling over policies, which is illustrated in Fig. 2. In stage I, a discovery policy $f$ is trained to figure out standing-up or rolling-over motions. $f$ is trained without deployment constraints, and only our task and symmetry rewards are used. In stage II, a deployable policy $\pi$ imitates the rolling-over / getting-up behaviors obtained from stage I under strong control regularization. This deployable policy $\pi$ is transferred from simulation to the real world as the final policy. To overcome the difficulty of contact-rich locomotion learning involved in these tasks, we design a curricula for learning. We detail the policy model and two-stage training in Section III-B, and the curriculum design is introduced in Section III-C.
@@ III-A Policy Architecture
HumanUP trains two policy models $f$ and $\pi$ with RL. Both policy models take observation $\mathbf{o}_{t}=[\mathbf{z}_{t},\mathbf{s}_{t},\mathbf{s}_{t-10:t-1}]\in\mathbb{R}^{868}$ as input and output action $\mathbf{a}_{t}\in\mathbb{R}^{23}$, where $\mathbf{s}_{t}\in\mathbb{R}^{74}$ is the robot’s proprioceptive information, $\mathbf{s}_{t-10:t-1}$ is the $10$ steps history states, and $\mathbf{z}_{t}\in\mathbb{R}^{54}$ are the encoded environment extrinsic latents learned using regularized online adaptation [18]. We use the proprioceptive information $\mathbf{s}_{t}$ consisting of the robot’s roll and pitch, angular velocity, DoF velocities, and DoF positions. Such proprioceptive information can be accurately obtained in the real world, and we find that these are sufficient for the robot to infer the overall posture. We do not use any linear velocity and yaw information as it is difficult to reliably estimate them in the real world [29].
The policy models are implemented as MLPs and trained via PPO [62]. The optimization targets maximizing the expected $\gamma$-discounted policy return within $T$ episode length: $\mathbb{E}\left[\sum_{t=1}^{T}\gamma^{t-1}r_{t}\right]$, where $r_{t}$ is the reward at timestamp $t$.
@@ III-B1 Stage I: Discovery Policy
This stage discovers getting-up / rolling-over behavior efficiently without deployment constraints. We use the following task rewards with very weak regularization to train this discovery policy $f$. We describe the rewards used for optimizing the two policies. Timestep $t$ and reward weight terms are omitted for simplicity. The precise expressions for each reward term and their weights are provided in supplementary.
@@
• $r_{\text{uprightness}}$ encourages the robot to increase the projected gravity on $z$ axis so that the robot will stand upright;
@@
• $r_{\text{symmetry}}$ reduces the search space by encouraging (but not requiring) the robot to output bilaterally symmetric actions. Past work [63, 33] employed hard symmetry which improves RL sample efficiency at the cost of limiting robots’ DoFs and generalization. Our soft symmetry reward inherits the benefit but mitigates the limitation.
Rewards for Rolling Over:  $r_{\text{roll}}=r_{\text{gravity}}$, where $r_{\text{gravity}}$ encourages the robot to change its body orientation so that its projected gravity is close to the target projected gravity when lying facing up.
@@ III-B2 Stage II: Deployable Policy
This stage trains policy $\pi$ that will be directly deployed in the real world. Policy $\pi$ is trained to imitate an 8$\times$ slowed-down version of the state trajectories discovered in Stage I, while also respecting strong regularization to ensure Sim2Real transferability. We use the typical control regularization rewards and describe them in supplementary. Below, we describe the tracking reward.
Tracking Rewards: The tracking reward, $r_{\text{tracking}}$, encourages the humanoid to act close to the given motion trajectory derived from the discovered motion. $r_{\text{tracking}}=r_{\text{tracking\_DoF}}+r_{\text{tracking\_body}}$, where
@@
• $r_{\text{tracking\_body}}$ encourages the robot to move the body to the same position as the reference. Specifically, $r_{\text{tracking\_body}}$ becomes $r_{\text{head\_height}}$ and $r_{\text{head\_gravity}}$ which encourage the robot to track the height and the projected gravity of the robot’s head for getting-up and rolling-over tasks, respectively.

### v2

@@ Abstract
Automatic fall recovery is a crucial prerequisite before humanoid robots can be reliably deployed. Hand-designing controllers for getting up is difficult because of the varied configurations a humanoid can end up in after a fall and the challenging terrains humanoid robots are expected to operate on. This paper develops a learning framework to produce controllers that enable humanoid robots to get up from varying configurations on varying terrains. Unlike previous successful applications of learning to humanoid locomotion, the getting-up task involves complex contact patterns (which necessitates accurately modeling of the collision geometry) and sparser rewards. We address these challenges through a two-phase approach that induces a curriculum. The first stage focuses on discovering a good getting-up trajectory under minimal constraints on smoothness or speed / torque limits. The second stage then refines the discovered motions into deployable (i.e. smooth and slow) motions that are robust to variations in initial configuration and terrains. We find these innovations enable a real-world G1 humanoid robot to get up from two main situations that we considered: a) lying face up and b) lying face down, both tested on flat, deformable, slippery surfaces and slopes (e.g., sloppy grass and snowfield). This is one of the first successful demonstrations of learned getting-up policies for human-sized humanoid robots in the real world.
Project page: https://humanoid-getup.github.io/
Figure 1: HumanUP provides a simple and general two-stage training method for humanoid getting-up tasks, which can be directly deployed on Unitree G1 humanoid robots [75]. Our policies showcase robust and smooth behavior that can get up from diverse lying postures (both supine and prone) on varied terrains such as grass slopes and stone tiles.
@@ I Introduction
This paper develops learned controllers that enable a humanoid robot to get up from varied fall configurations on varied terrains. Humanoid robots are susceptible to falls, and their reliance on humans for fall recovery hinders their deployment. Furthermore, as humanoid robots are expected to work in environments involving complex terrains and tight workspaces (i.e. challenging scenarios that are too difficult for wheeled robots), a humanoid robot may end up in an unpredictable configuration upon a fall, or may be on an unknown terrain. 26 of the 46 trials at the DARPA Robotics Challenge (DRC) had a fall, and 25 of these falls required human intervention for recovery [44]. The DRC identified fall prevention and recovery as a major topic needing more research. This paper pursues it and proposes a learning-based framework for learning fall recovery policies for humanoid robots under varying conditions.
@@
We present HumanUP, a two-stage reinforcement learning (RL) training framework that circumvents these issues. Stage I targets solving the task in easier settings (sparse task rewards with weak regularization), while Stage II makes the learned motion deployable (i.e., control should be smooth; velocities and executed torques should be small; etc). Discovering the getting-up motion is hard because of sparse and underspecified rewards. Stage I tackles this hard problem without being limited by smoothness in motion or speed / torque limits. Tracking a trajectory is easier as it offers dense rewards. Stage II tackles this easier problem but does it under strict Sim2Real control regularization and randomization of terrains and initial poses. Thus, going from Stage I to Stage II corresponds to a learning curriculum that progresses from simplified $\rightarrow$ full collision mesh, canonical $\rightarrow$ random initial lying posture, and weak $\rightarrow$ strong control regularization, and domain randomization. This amounts to a hard-to-easy curriculum on task difficulty (Stage I: getting-up task; Stage II: motion tracking), and an easy-to-hard curriculum on regularization and variability (Stage I: weaker, Stage II: stronger).
We conduct experiments in simulation and the real world with the G1 platform from Unitree. In the real world, we find our framework enables the G1 robot to get up from two different poses (supine, i.e. lying face up, and prone, i.e. lying face down) across six different terrains. This expands the capability of the G1 robot: the manufacturer-provided hand-crafted getting-up controller only successfully gets up from supine poses on a flat surface without bumps. In simulated experiments, our framework can successfully learn getting-up policies that work on varied terrains and varied starting poses.
@@ II-A Humanoid Control
Controlling a high degree of freedom humanoid robots has fascinated researchers for the last several decades. Model-based techniques, such as those based on the Zero Moment Point (ZMP) principle [76, 35, 65, 74], optimization [45, 4, 14], and Model Predictive Control (MPC) [12, 22, 15, 79], have demonstrated remarkable success in fundamental locomotion tasks like walking, running and jumping. However, these approaches often struggle to generalize or adapt to novel environments. In contrast, learning-based approaches have recently made significant strides, continuously expanding the generalization capabilities of humanoid locomotion controllers.
@@ II-A1 Learning for humanoid control
Figure 2: HumanUP system overview. Our getting-up policy (Sec. III-A) is trained in simulation using two-stage RL training, after which it is directly deployed in the real world. (a) Stage I (Sec. III-B1) learns a discovery policy $f$ that figures out a getting-up trajectory with minimal deployment constraints. (b) Stage II (Sec. III-B2) converts the trajectory discovered by Stage I into a policy $\pi$ that is deployable, robust, and generalizable. This policy $\pi$ is trained by learning to track a slowed down version of the discovered trajectory under strong control regularization on varied terrains and from varied initial poses. (c) The two-stage training induces a curriculum (Sec. III-C). Stage I targets motion discovery in easier settings (simpler collision geometry, same starting poses, weak regularization, no variations in terrain), while Stage II solves the task of making the learned motion deployable and generalizable.
@@ II-B1 Getting up via motion planning
Early work from Morimoto and Doya [58] solved the getting-up problem for a two-joint, three-link walking robot in 2D, and several discrete states are used as subgoals to transit via hierarchical RL. This line of work can be viewed as an application of motion planning by configuration graph transition learning [43], where stored robot states between lying and standing are used as graph nodes to transit [21, 40, 41, 69]. More recently, some progress has been made to enable toy-sized humanoid robots to get up [26, 37, 67, 24]. For example, González-Fierro et al. [26] explores getting up from a canonical sitting posture with motion planning by imitating human demonstration with ZMP criterion. To address the high-dimensionality of humanoid configurations, Jeong and Lee [37] leverage bilateral symmetry to reduce the control DoFs by half and a clustering technique is used for further reducing the complexity of configuration space, thereby improving getting-up learning efficiency. However, such state machine learning using predefined configuration graphs may not be sufficient for generalizing to unpredictable initial and intermediate states, which happens when the robot operates on challenging terrains.
@@ II-B2 Hand-designed getting-up trajectories
Another solution, often adopted by commercial products, is to replay a manually designed motion trajectory. For example, Unitree [75] has a getting-up controller built into G1’s default controllers. Booster Robotics [1] designed a specific recovery controller for their robots that can help the robot recover from fallen states. Concurrent work from Zhuang and Zhao [82] enables a G1 robot to get up by tracking the getting-up motion of a real human. The main drawback of such pre-defined trajectory getting-up controllers is that they may only handle a limited number of fallen states and lack generalization.
@@ II-B3 Learned getting-up policies for real robots
RL followed by sim-to-real has also been successfully applied for quadruped [47, 39, 77, 55] fall recovery. For example, Lee et al. [47] explore sim2real RL to achieve real-world quadruped fall recovery from complex configurations. Ji et al. [39] train a recovery policy that enables the quadruped to dribble in snowy and rough terrains continuously. Wang et al. [77] develop a quadruped recovery policy in highly dynamic scenarios.
@@ II-B4 Learned getting-up policies for character animation
A parallel research effort in character animation, also explores the design of RL-based motion imitation algorithms: DeepMimic [59], AMP [60], PHC [53], among others [11, 54, 72, 5, 23, 78]. These have also demonstrated successful getting-up controllers in simulation. By tracking user-specified getting-up curves, Frezzato et al. [17] enable humanoid characters to get up by synthesizing physically plausible motion. Without recourse to mocap data, such naturalistic getting-up controllers for simulated humanoid characters can also be developed with careful curriculum designs [70]. Some works explore sampling-based methods for addressing contact-rich character locomotion, including getting up [49, 31, 61], while some works have demonstrated success in humanoid getting up with online model-predictive control [71]. It is worth noticing, however, that these works use humanoid characters with larger DoFs compared to humanoid robots (e.g., 69 DoFs in SMPL [51]) and use simplified dynamics. As a result, learned policies operate body parts at high velocities and in infeasible ways, leading to behavior that cannot be transferred into the real world directly. Hence, developing generalizable recovery controllers for humanoid robots remains an open problem.
@@ III HumanUP: Sim-to-Real Humanoid Getting Up
Our goal is to learn a getting-up policy $\pi$ that enables a humanoid to get up from arbitrary initial postures. We consider getting up from two families of lying postures: a) supine poses (i.e. lying face up) and b) prone poses (i.e. lying face down). Getting up from these two groups of postures may require different behaviors, which makes it challenging to learn a single policy that handles both scenarios. To tackle this issue, we decompose the getting-up task from a prone pose to first rolling over and then standing up from the resulting supine posture. Therefore, we aim to learn policies for rolling over from a prone pose and getting up from a supine pose separately.
To solve these two tasks, we propose HumanUP, a general learning framework for training getting-up and rolling over policies, which is illustrated in Fig. 2. In Stage I, a discovery policy $f$ is trained to figure out standing-up or rolling-over motions. $f$ is trained without deployment constraints, and only the task and symmetry rewards are used. In stage II, a deployable policy $\pi$ imitates the rolling-over / getting-up behaviors obtained from stage I under strong control regularization. This deployable policy $\pi$ is transferred from simulation to the real world as the final policy. We detail the policy model and two-stage training in Sec. III-B, and then discuss the induced curriculum in Sec. III-C.
@@ III-A Policy Architecture
HumanUP trains two policy models $f$ and $\pi$ with RL. Both policy models take observation $\mathbf{o}_{t}=[\mathbf{z}_{t},\mathbf{s}_{t},\mathbf{s}_{t-10:t-1}]\in\mathbb{R}^{868}$ as input and output action $\mathbf{a}_{t}\in\mathbb{R}^{23}$, where $\mathbf{s}_{t}\in\mathbb{R}^{74}$ is the proprioceptive information, $\mathbf{s}_{t-10:t-1}$ is the $10$ steps history states. $\mathbf{z}_{t}\in\mathbb{R}^{54}$ are the encoded environment extrinsic latents that are predicted from observation history and learned using regularized online adaptation [19]. The proprioceptive information $\mathbf{s}_{t}$ consists of the robot’s roll and pitch, angular velocity, DoF velocities, and DoF positions. Such proprioceptive information can be accurately obtained in the real world, and we find that this is sufficient for the robot to infer the overall posture. We do not use any linear velocity and yaw information as it is difficult to reliably estimate them in the real world [33, 32].
The policy models are implemented as MLPs and trained via PPO [66]. The optimization maximizes the expected $\gamma$-discounted policy return within $T$ episode length: $\mathbb{E}\left[\sum_{t=1}^{T}\gamma^{t-1}r_{t}\right]$, where $r_{t}$ is the reward at timestamp $t$.
@@ III-B1 Stage I: Discovery Policy
This stage discovers getting-up / rolling-over behavior efficiently without deployment constraints. We use the following task rewards with very weak regularization to train this discovery policy $f$. Timestep $t$ and reward weight terms are omitted for simplicity. The precise expressions for each reward term and their weights are provided in Sec. A.1.
@@
• $r_{\text{uprightness}}$ encourages the robot to increase the z-component of the projected gravity, [^1] Projected gravity on a robot part is the gravity vector transformed from the world frame to the part’s local frame. so that the robot stands upright;
@@
• $r_{\text{symmetry}}$ reduces the search space by encouraging (but not requiring) the robot to output bilaterally symmetric actions. Past work [68, 37] employed hard symmetry which improves RL sample efficiency at the cost of limiting robots’ DoFs and generalization. Our soft symmetry reward partially leverages the benefit but mitigates the limitation.
Rewards for Rolling Over:  $r_{\text{roll}}=r_{\text{gravity}}$, which encourages the robot to change its body orientation so that its projected gravity is close to the projected gravity when lying face up.
@@ III-B2 Stage II: Deployable Policy
This stage trains policy $\pi$ that will be directly deployed in the real world. Policy $\pi$ is trained to imitate an 8×8$\times$8 × slowed-down version of the state trajectories discovered in Stage I, while also respecting strong regularization to ensure Sim2Real transferability. We use the typical regularization rewards and describe them in Sec. A.2. Below, we describe the tracking reward.
Tracking Rewards: $r_{\text{tracking}}$ encourages the robot to act close to the given motion trajectory derived from the discovered motion. $r_{\text{tracking}}=r_{\text{tracking\_DoF}}+r_{\text{tracking\_body}}$, where
@@
• $r_{\text{tracking\_body}}$ encourages the robot to move the body to the same position as the reference. Specifically, $r_{\text{tracking\_body}}$ becomes two different rewards to encourage tracking upright posture ($r_{\text{head\_height}}$) and correct head orientation ($r_{\text{head\_gravity}}$) for getting-up and rolling-over tasks, respectively.